OpenStudy (anonymous):

Solve the equation by finding square roots 32x^2=50

OpenStudy (anonymous):

start by diving each side by 32

OpenStudy (anonymous):

I have and got 2 now i have no idea what to do. The answers are -4/5,4/5 -25/16, 25/16 -16/25, 16/25 -5/4, 5/4

OpenStudy (zehanz):

Try to write everything down step by step. Begin with surdawi's idea:\[32x^2=50 \rightarrow x^2=\frac{ 32 }{ 50 }\]You could now go on and take the root straightwawy, but it is much better to simplify the fraction first!\[x^2=\frac{ 32 }{ 50 }=\frac{ 16 }{ 25 }=\frac{ 4^2 }{ 5^2 }\]Now I think you can see the two numbers that satisfy the equation!

OpenStudy (zehanz):

Sorry, typo! I did it upside down :( It is of course:\[x^2=\frac{ 50 }{ 32 }=\frac{ 25 }{ 16 }=\frac{ 5^2 }{ 4^2 }\]

OpenStudy (anonymous):

so it is -5/4, 5/4 ?

OpenStudy (zehanz):

It is!!

OpenStudy (anonymous):

would you mind helping me with another one?

OpenStudy (zehanz):

No problem!